OpenStudy (anonymous):

What is the value of the y variable in the solution to the following system of equations? 3x − 6y = 3 7x − 5y = −11

OpenStudy (anonymous):

Use substitution.

OpenStudy (anonymous):

._. ... 0 4 there is no value, or x can be any number?

OpenStudy (anonymous):

Here's what I mean: \(3x-6y=3 \rightarrow 3x=3+6y \rightarrow x=1+2y\). Put that value of \(x\) in terms of \(y\) into your second equation and solve for \(y\). Then it becomes easy to find \(x\).

OpenStudy (anonymous):

._. nvm dude.

OpenStudy (anonymous):

3x − 6y = 3 7x − 5y = −11 3x=3+6y x=1+2y 7(1+2y)-5y=-11 7+14y-5y=-11 7+9y=-11 9y=-11-7 =-18 y=-18/9 =-2 7x-5(-2)=-11 7x+10=-11 7x=-11-10 7x=-21 x=-21/3 x=-7 Hope This Helped :)

OpenStudy (anonymous):

thanks o:

OpenStudy (anonymous):

:) no prob